Please help me solve this problem: 2(x+4)=5x-3
Answer by Alan3354(69443)

Please help me solve this problem: 2(x+4)=5x-3
-----------
That's not a problem, it's an equation.
---
The only variable is x, so I think you want to solve for x.
==========
You should say, "Solve for x."
But, at least you didn't say, "What is... ?"
-------------
2(x+4) = 5x-3
2x+8 = 5x - 3
Add 3
2x + 11 = 5x
Subtract 2x
11 = 3x
Divide by 3
11/3 = x
x = 11/3